For answering Questions 17 to 20, refer to the following figures extracted from the financial statements of Blunt Ltd. Sales Cost of sales Profit before interest and tax Fixed assets Depreciation Stock Debtors Cash at bank Creditors Long-term loans 2,700 1,950 300 1,500 150 1,050 750 150 900 450 17) What is the quick ratio? a) 100-1 b) 217:1 c) 083-1 d) 333:1 e) none of the above 18) What is the net working capital? a) 1,050 b) 1,950 c) 900 e) none of the above 19) What is the debtor collection period in days, assuming a 365-day year? a) 63 days b) 101 days c) 140 days d) 122 days e) none of the above 20) What is the gross profit margin? a) b) c) d) e) 72.2% 11.1% 27.7% 38.5% none of the above
Answer #1

17)

Quick ratio = ( cash + marketable securities + debtors) / Current liabilities

Quick ratio = ( 150 + 750) / 900

Quick ratio = 1.00:1

18)

Net working capital = Current assets - current liablities

Net working capital = Cash + debtors + stock - creditors

Net working capital = 150 + 750 + 1,050 - 900

Net working capital = 1,050

19)

Debt turnover ratio = sales / receivables

Debt turnover ratio = 2,700 / 750

Debt turnover ratio = 3.6

Debt collection period = 365 / 3.6

Debt collection period = 101 days

20)

Gross profit = 2,700 - 1,950 = 750

Gross profit margin = ( gross profit / sales ) * 100

Gross profit margin = ( 750 / 2,700) * 100

Gross profit margin = 27.7%